TOGO Estimate the H-F bond enthalpy knowing that the Heat of Formation of hydrogen fluoride (AH;, HF) is -271 kJ/mol.
Answer #1

The reaction is

H2 + F2\rightarrow 2 HF (1)

Bond Bond enthalpy, KJ/mol
H-H 436
F-F 157
H-F unknown
Substance Heat of formation (KJ/mole)
H2 0
F2 0
H-F -271

From reaction 1.

Heat of formation of 2× HF = \sum Bond enthalpy of reactants - \sumBond enthalpy of products

Or, -271× 2 = (436+157) - 2× Bond enthalpy of HF

Or, 2× Bond enthalpy of HF = - 271×2 - (436 +157)

or, 2× Bond enthalpy of HF = - 1135.3 KJ

Or, Bond enthalpy of HF = (1135 /2) = 567.5 KJ/mole.